OpenStudy (anonymous):

Estimate the value of each square root to the tenths place: 39 I've have no idea on how to even start.. please help?

jimthompson5910 (jim_thompson5910):

you can use a calculator to evaluate sqrt(39) or you can do so by hand (check out the page below) http://www.homeschoolmath.net/teaching/square-root-algorithm.php

OpenStudy (anonymous):

Okay so I got the two consecutive numbers and I got 36 and 49 then I said there was 13 in between so then it would have to be 13. Something but whenever I multliply it by itself it comes out WAY over what it's suppose to be close to. Sorry if I don't make any sense but I'm really trying

jimthompson5910 (jim_thompson5910):

since 6^2 = 36 and 7^2 = 49 and because 39 is between 36 and 49, this means that sqrt(39) must be between 6 and 7

OpenStudy (anonymous):

What?

jimthompson5910 (jim_thompson5910):

sqrt(39) means "square root of 39"

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

After the consecutive part because I thought you just counted how many were in between them and started guessing...

jimthompson5910 (jim_thompson5910):

you could do that too we know for a fact that sqrt(39) is between 6 and 7 so start guessing, so guess 6.1, 6.2, etc and check each guess by squaring it

OpenStudy (anonymous):

So did I get the consecutive numbers wronged? Because I got them as 36 & 49 so I got 13 in between that's where I'm confused.. I'm so sorry...

jimthompson5910 (jim_thompson5910):

how are you getting 13?

jimthompson5910 (jim_thompson5910):

I see how you're getting 36 and 49

OpenStudy (anonymous):

I thought my math teacher said something about counting how many numbers are between 36 & 49 so I counted and got 13.

jimthompson5910 (jim_thompson5910):

oh i see

jimthompson5910 (jim_thompson5910):

I'm not sure why she told you to do that

jimthompson5910 (jim_thompson5910):

notice how 6^2 = 36 so that means the square root of 36, sqrt(36) for short, is 6 ie sqrt(36) = 6 make sense so far?

OpenStudy (anonymous):

Yep

jimthompson5910 (jim_thompson5910):

and also notice how 7^2 = 49 so sqrt(49) = 7

jimthompson5910 (jim_thompson5910):

because sqrt(36) = 6 and because sqrt(49) = 7 this means that sqrt(39) must be between 6 and 7

jimthompson5910 (jim_thompson5910):

since 39 is between 36 and 49

OpenStudy (anonymous):

Gotcha!

jimthompson5910 (jim_thompson5910):

so you just have to go through 6.1, 6.2, 6.3, 6.4, etc and square each one and see which one gets closest to 39

OpenStudy (anonymous):

I've got no clue what I was doing but you helped me A LOT. So my answer should be 6.2 because that's the closest I can get to 39 without going over right?

jimthompson5910 (jim_thompson5910):

yes it turns out (by use of a calculator) that sqrt(39) = 6.244997998 and this is approximate of course

jimthompson5910 (jim_thompson5910):

so 6.2 is as close as you get

OpenStudy (anonymous):

Yep thanks (:

jimthompson5910 (jim_thompson5910):

you're welcome